OpenStudy (johan14th):

can someone help me find the power form of the reciprocal radical form 1/(3 sqrt.three(x^2))

OpenStudy (johan14th):

\[\frac{ 1 }{ 3\sqrt[3]{x ^{2}}}\]

OpenStudy (anonymous):

you have a simple formula for radicals \[\sqrt[a]{x^b}=x^\frac{b}{a}\]